OpenStudy (anonymous):

@hero Find the unit rate, and then use it to solve the problem. A trucking company makes 50 deliveries in 5 days. How long will it take for 400 deliveries? A. 8 days B. 10 days C. 20 days D. 40 days

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what the unit rate?

OpenStudy (amistre64):

think, how long does it take to make 1 (a single unit) delivery based on the given information? 1 delivery per ?? days

OpenStudy (anonymous):

i don't get this question at all

OpenStudy (amistre64):

i just explained it ....

OpenStudy (amistre64):

if we make 50 deliveries in 5 days; then for some scalar value, n, we can make 50n deliveries in 5n days so the question is: when does 50n = 1? OR to reduce the computations .... when does 50n = 400?

OpenStudy (anonymous):

8

OpenStudy (amistre64):

50n = 400 when n=8 yes :) but thats not the unit rate, its the straight forward solution so, it takes us 5n days to do it ... how many days then?

OpenStudy (anonymous):

C

OpenStudy (amistre64):

C is 100 in roman numerals so im going to say no ... it takes us 5*8 = 40 days ;)

OpenStudy (anonymous):

An automated car wash can wash 30 cars in 5 hours. How long would it take to wash 120 cars? A. 15 h B. 20 h C. 25 h D. 30 h

OpenStudy (amistre64):

scale the ratio ... what do we get?

OpenStudy (amistre64):

if you know how to work simliar triangles or rectangles ... thats all this is

OpenStudy (amistre64):



OpenStudy (anonymous):

30c=120

OpenStudy (amistre64):

good so far, i like it

OpenStudy (anonymous):

B

OpenStudy (amistre64):

yep, the scalar is 4, so 5*4 = 20

OpenStudy (anonymous):

A worker at a computer factory can assemble 8 computers per hour. How long would it take this worker to assemble 200 computers? A. 8 h B. 20 h C. 25 h D. 200 h

OpenStudy (amistre64):

8 in 1 hour 8n in n hours ... whats the scalar?

OpenStudy (amistre64):

if anyting, 80 in 10 160 in 20 .. so a little more than 20 seems logical

OpenStudy (anonymous):

B.

OpenStudy (amistre64):

not B .... 200 computers not 160 ...

OpenStudy (anonymous):

D

OpenStudy (amistre64):

do the math :/

OpenStudy (anonymous):

am i still wrong

OpenStudy (amistre64):

you tell me, chk your results does 8*200 = 200?

OpenStudy (amistre64):

does 8*20 = 200? does 8*8 = 200? does 8*25 = 200?

OpenStudy (anonymous):

8x25=200

OpenStudy (anonymous):

C